# Delay-Aware Robust Edge Network Hardening Under Decision-Dependent Uncertainty

### v1

@@ II-C Uncertainty Modeling
RO-DIU: $\min_{{\bm{t}},{\bm{x}},{\bm{w}}}~{}\max_{{\bm{d}}\in\mathcal{D}_{1}}\>\mathcal{C}_{1}+\mathcal{C}_{2}$ (13a)
@@
RO-NH: $\min_{{\bm{x}}\in\mathbb{Z}_{+}^{I\times J},{\bm{w}}\in\mathbb{Z}_{+}^{I}}\max_{{\bm{d}}\in\mathcal{D}_{1}}\>\mathcal{C}_{2}$ (14)
@@
2. Robust Hardening Model with Endogenous Uncertainties: Link hardening can enhance the quality of the edge network by reducing link delay and delay variation. However, the link delay tends to increase as data traffic (workload) traversing through the link increases. Hence, the set $\mathcal{D}_{1}(\hat{d},\bar{d},\Gamma_{1})$ fails to capture the interdependencies between the actual link delays and link hardening as well as workload allocation decisions. To overcome this limitation, we propose a new uncertainty set that incorporates decision-dependent uncertainty to accurately model the impact of decision variables on the actual link delays. This approach enables us to better capture the effects of link hardening and workload allocation decisions on the actual link delays, thereby improving the model’s robustness and efficiency. The endogenous uncertainty set representing the link delay uncertainty is defined as follows:
@@
Here, we introduce two new decision-dependent parameters, $\gamma_{i,j}^{r}$ and $u_{i,j}$. The parameter $\gamma_{i,j}^{r}$ captures the impact of link hardening on the network delay along the link $(i,j)$ with level-$r$ hardening. As the level of hardening increases, the link delay deviation decreases. Note that at most one hardening level can be chosen for each link, as indicated in equation (5).
The network delay on a link depends on the link utilization, which is in turn influenced by the amount of data traffic or workload on that link. For simplicity, we assume that the network delay on each link is a linear function of its workload, represented by the slope parameter $u_{i,j}$. This means that the link delay will increase proportionally with an increase in workload. Note that the dependence of the link delay on the workload allocation decision can be represented using more complex functions such as piecewise linear functions.
@@
RODDU: $\min_{{\bm{t}},{\bm{x}},{\bm{w}}}~{}\max_{{\bm{d}}\in\mathcal{D}_{2}}\>\mathcal{C}_{1}+\mathcal{C}_{2}$ (16a)
@@ III Solution Approach
This section develops exact reformulations to solve the robust model (16). The presence of decision variables, which are solutions to the outer minimization problem, in the uncertainty set $\mathcal{D}_{2}$ poses significant complications to solving RODDU. Specifically, it changes the size and form of the set, making it difficult to reformulate robust problems with DDU sets in a tractable fashion. Indeed, the general RODDU problem is NP-complete, which can be proved based on the 3-Satisfiability problem (i.e., 3-SAT) [18]. Please refer to Appendix. -A for the NP-Complete proof. Moreover, the uncertain parameter $d_{i,j}$ appears in both the objective function and the left-hand side of (16c), and the problem is a nonlinear program containing bilinear terms, such as $d_{i,j}x_{i,j}$, that cannot be solved directly. To overcome these challenges, we present two solution approaches. The first approach combines linear programming (LP) duality, binary expansion, and linearization methods to reformulate the RODDU model into a large-scale MILP that can be solved using solvers. The second approach introduces new auxiliary variables and a clever transformation trick to significantly reduce the size of the problem by removing redundant constraints during the problem reformulation, which can substantially accelerate the computational time.
III-A Robust Solution
The main challenge in solving the RODDU problem stems from the fact that the objective function (16a) and the robust constraints (16c) contain bilinear terms due to the inclusion of decision variables in the DDU set $\mathcal{D}_{2}$. When the uncertainty set is independent of the decision variables, standard techniques can be employed to convert the problem into an equivalent MILP robust counterpart using LP duality [15]. Unfortunately, $\mathcal{D}_{2}$ is decision-dependent, rendering our problem more complicated. To this end, we introduce a series of transformations to convert RODDU into an equivalent and solvable MILP form by transforming (16a) and (16c) into a set of linear equations.
@@ III-A Robust Solution
By using the preceding linearization steps, we obtain the following linear constraints that are equivalent to (20a)-(20c).
@@
The last term in (III-A) can be written explicitly as:
@@
Finally, the RODDU problem (16) can be reformulated as the following MILP, which can be solved by off-the-shelf solvers.
@@
[…]
III-B Enhanced Robust Solution
Although the MILP reformulation (30) can solve the decision-dependent robust problem (RODDU) in (16), it suffers from weak relaxations due to the big-M terms in (21), leading to poor numerical performance. Furthermore, the size of the reformulation RDDU increases rapidly as the problem size, especially the upper bound of $x_{i,j}$, increases. To this end, we propose an alternative reformulation that can substantially improve the computational time. Specifically, we introduce new auxiliary variables and employ an alternative but equivalent form of $\mathcal{D}_{2}$ to reformulate the original RODDU problem. We first rewrite the constraints on $g_{i,j}$ in $\mathcal{D}_{2}$ as follows:
@@ III-B Enhanced Robust Solution
$\mathcal{D}_{2}(\hat{d},\bar{d},t,x,\Gamma_{2}):=\bigg{\{}d:~{}g\in\mathcal{G}^{\prime},$
$\!d_{i,j}\!=\!\bar{d}_{i,j}\!+\!\hat{d}_{i,j}\big{(}g_{i,j}+u_{i,j}x_{i,j}\big{)},\forall i,j\bigg{\}},$ (32)
@@
It can be observed that the key difference between the enhanced robust solution and the robust solution in Section III-A lies in the introduction of new binary variables ($v_{i,j}^{r}=1-t_{i,j}^{r}$) and replacing the upper bound of $g$ in (II-C) with (31). Now, the robust delay constraint (16c) can be expressed as:
@@
The last term in (III-B) can be given explicitly as:
@@
Overall, the RODDU problem (16) can be reformulated as the following equivalent MILP:
@@
[…]
III-C Equivalence of RDDU and e-RDDU
@@
$\mathcal{U}({\bm{z}}):=\bigg{\{}\bm{\zeta}|\bm{A}\bm{\zeta}\leq{\bm{v}}+\bm{\psi}{\bm{z}}\bigg{\}},$ (41)
@@
It is easy to see that the RDDU reformulation relies on the result of Theorem III.1, while e-RDDU exploits Theorem III.2 to reduce the number of Big-M constraints. Therefore, we can infer that RDDU and e-RDDU are equivalent. Also, the computational advantages of e-RDDU over RDDU stem from the fact that (44) has considerably fewer integer constraints compared to (43).
III-D Comparison Between RDDU and e-RDDU
@@ IV-A Simulation Setting
We consider an EC system comprising 10 areas and 10 ENs (i.e., $I$ = $J$ = $10$). We will also examine larger networks in our sensitivity analyses. The edge network topology is based on the cities and locations of randomly selected Equinix edge data centers [^3] https://www.equinix.com/data-centers/americas-colocation. The network delay parameters, including the minimum delay $\bar{d}_{i,j}$ and maximum delay deviation $\hat{d}_{i,j}$, are generated using the global ping dataset444 https://wondernetwork.com/pings. By exploiting the Materna data trace [^5] http://gwa.ewi.tudelft.nl/datasets/gwa-t-13-materna, we randomly generate resource demands $\lambda_{i}$’s following a uniform distribution with values ranging from 40 to 60 vCPUs. We consider three hardening levels for each link (i.e., $R=3$). The link hardening cost parameters (i.e., $h_{i,j}^{r},\forall i,j,r$) are generated as follows. Recall that $\Delta h$ is denoted by the incremental link hardening cost between two adjacent levels for each link, which is set to 0.2 in our simulation. The first-level hardening costs $h_{i,j}^{\sf 1}$ are randomly generated from a uniform distribution within the interval $[1,1.05]$. The second and third-level link hardening costs are obtained by adding the incremental cost to the first-level hardening costs, i.e., $h_{i,j}^{\sf 2}=h_{i,j}^{\sf 1}+\Delta h$ and $h_{i,j}^{\sf 2}=h_{i,j}^{\sf 1}+2\Delta h,~{}\forall i,j$.
The parameters $\gamma_{i,j}^{r}$, which represent the impact of the hardening decision on the link delay, are generated using a similar method. Specifically, we set $\gamma_{i,j}^{1},\forall i,j$ to 0.1 and define $\Delta\gamma$ as an incremental impact factor between adjacent levels, which is set to 0.4. Then, we have $\gamma^{2}=\gamma^{1}+\Delta\gamma$ and $\gamma^{3}=\gamma^{1}+2\Delta\gamma$. For simplicity, we assume that the workload has the same impact on every link. Thus, in the base case, $u_{i,j}={\bm{u}}=0.1,\forall i,j$. The resource capacities of ENs are randomly selected based on EC2 instances [^6] https://instances.vantage.sh.. The unmet demand penalty parameters $s_{i}$ are generated following a uniform distribution on $[40,50]$. In our default setting, the values of other system parameters are as follows: $\rho=0.1$, $\alpha=\alpha_{i}=0.05,~{}\forall i$, $\Delta\gamma=0.4$, $B=100$, $\Gamma_{1}=\Gamma_{2}=\Gamma=15$, ${\bm{u}}=0.2$, and $\Delta_{i}=\Delta=15$. During our sensitivity analysis, we will also vary these important system parameters. All the experiments are implemented in a Matlab environment using CVX [^7] https://cvxr.com/cvx/ and Gurobi 9.1.2 on a desktop with an Intel Core i7-11700KF and 32 GB of RAM.
@@ IV-B1 Impacts of the link hardening costs and budget
As illustrated in Figs.1(a) and 1(b), increasing the hardening budget B leads to a decrease in the total cost since the platform can select more links for hardening, resulting in a reduction of the total cost. It is important to note that the delay constraints (9) prioritize the closest ENs to meet the demand of each area, leading to only a subset of logical links being utilized. Clearly, the platform should only harden links with traffic traversing through them. Thus, as shown in Fig. 1(a), the cost becomes saturated after a certain budget value since the platform has already chosen the most critical links for hardening and has no incentive to harden more links even if the budget is redundant.
@@ IV-B2 Impacts of the uncertainty set
Figs. 2(a)-2(b) exhibit the effects of the uncertainty set on the optimal solution. It is evident from Fig. 2(a) that the total cost initially increases with the increment of $\Gamma$, resulting in a larger uncertainty set and a more robust solution. However, the total cost reaches saturation after a certain value of $\Gamma$, as the platform has already chosen an optimal set of links for hardening. Note that only a subset of ENs can serve demand from a specific area, causing certain logical links to remain unused for data traffic delivery. Therefore, the platform has no incentive to harden these links. Recall that $\bf{u}$ represents the influence of workload on link delay, as specified in $\mathcal{D}_{2}$. Fig. 2(b) suggests that a higher value of $\bf{u}$ corresponds to a greater link delay variation, leading to a higher total cost. These results are intuitive, as a larger uncertainty set yields a more robust solution, which, in turn, raises the cost.
@@
Figs. 3(a)-3(d) demonstrate the impacts of link hardening cost parameters (i.e., $\Psi$) and link hardening impact factors (i.e., $\gamma$) on the optimal objective and hardening cost (i.e., the total payment for link hardening). The incremental impact ($\Delta\gamma$) between two adjacent link hardening levels is also considered, where a higher value of $\Delta\gamma$ indicates a greater impact of level-2 and level-3 hardening, thereby offering more incentives for the platform to select links for level-2 and level-3 hardening. From $\mathcal{D}_{2}$, an increase in the link hardening impact results in lower link delays, leading to a reduction in overall delay cost. As $\Delta\gamma$ increases, so does its impact.
Fig. 3(a) shows that as $\Delta\gamma$ increases, the total cost decreases due to the significant reduction in link delay achieved through link hardening. This finding is further supported by Figs. 3(e)3(f). In addition, Fig. 3(b) indicates that the link hardening payment initially increases with the link hardening cost scaling factor $\Psi$ and then decreases after a certain value of $\Psi$, beyond which the hardening cost outweighs the benefit.
In Figs. 3(c)3(d), we set the incremental impact factor $\Delta\gamma$ to 0.3 and vary first-level impact factor ($\bm{\gamma}^{\sf 1}=\gamma_{i,j}^{\sf 1},\forall i,j$) from $0.1$ to $0.4$. A larger value of $\gamma^{\sf 1}$ leads to a lower link delay. Fig. 3(c) reveals that the total cost decreases as $\gamma^{\sf 1}$ increases, while Fig. 3(d) implies the platform invests more in link hardening as $\gamma^{\sf 1}$ increases. Finally, we vary $\Delta\gamma$ and the uncertainty budget $\Gamma$ to study their impact on the total cost and the link hardening payment in Figs. 3(e)3(f). It can be seen that an increase in $\Gamma$ leads to a higher cost due to the larger uncertainty set. Moreover, Fig. 3(f) suggests that the platform is willing to invest more money in link hardening as $\Gamma$ increases, indicating a higher level of uncertainty.

### v2

@@ II-B Deterministic Hardening Model
Remark II.2.
The constraint (1c) restricts each link $(i,j)$ to at most a single hardening level, which simplifies the model by preclassifying and grouping hardening technologies into distinct levels. Higher levels may incorporate the protections of lower levels, thereby eliminating the need to select multiple levels simultaneously. In contrast, the alternative model of selecting multiple levels to combine techniques may require additional constraints and complicate the formulation, especially when certain combinations are infeasible or incompatible.
@@ II-C Uncertainty Modeling
(RO-DIU) $\min_{{\bm{t}},{\bm{x}},{\bm{w}}}~{}\max_{{\bm{d}}\in\mathcal{D}_{1}}~{}~{}\mathcal{C}_{1}+\mathcal{C}_{2}$ (4a)
@@
(RO-NH) $\min_{{\bm{x}}\in\mathbb{Z}_{+}^{I\times J},{\bm{w}}\in\mathbb{Z}_{+}^{I}}\max_{{\bm{d}}\in\mathcal{D}_{1}}\>\mathcal{C}_{2}$ (5)
@@
2. Robust Hardening Model with Endogenous Uncertainties: Link hardening can enhance the quality of the edge network by reducing link delay and delay variation. However, the link delay tends to increase as data traffic (workload) traversing through the link increases. Hence, the set $\mathcal{D}_{1}(\hat{d},\bar{d},\Gamma_{1})$ fails to capture the interdependencies between the actual link delays and link hardening as well as workload allocation decisions. To overcome this limitation, we propose a new uncertainty set that incorporates DDU to accurately model the impact of decision variables on the actual link delays. This approach enables us to better capture the effects of link hardening and workload allocation decisions on the actual link delays, thereby improving the model’s robustness and efficiency. The endogenous uncertainty set representing the link delay uncertainty is defined as follows:
@@
Here, we introduce two new decision-dependent parameters, $\gamma_{i,j}^{r}$ and $u_{i,j}$. The parameter $\gamma_{i,j}^{r}$ captures the impact of link hardening on the network delay along the link $(i,j)$ with level-$r$ hardening. As the level of hardening increases, the link delay deviation decreases. Note that at most one hardening level can be chosen for each link, as indicated in equation (1c).
Remark II.3 (On Linear Approximation of Network Delay).
The network delay on a given link is influenced by its utilization, which is determined by the volume of data traffic or workload on that link. For simplicity, in the endogenous uncertainty set described in (II-C), the network delay on each link is modeled as a linear function of its workload, parameterized by the slope coefficient $u_{i,j}$. This linear approximation is justified in scenarios where the network operates within capacity constraints and congestion is effectively managed [18].
However, this assumption has inherent limitations. In scenarios with high traffic volumes or significant congestion, the relationship between latency and workload may exhibit non-linear characteristics due to various factors such as queuing delays and packet loss. To better capture these complexities, more advanced modeling approaches or linearization techniques, such as piecewise linear approximations, can be employed. The proposed model provides a foundational framework that can be extended to accommodate more complex scenarios.
@@
(RO-DDU) $\min_{{\bm{t}},{\bm{x}},{\bm{w}}}~{}\max_{{\bm{d}}\in\mathcal{D}_{2}}~{}~{}\mathcal{C}_{1}+\mathcal{C}_{2}$ (7a)
@@ III Solution Approach
This section develops exact reformulations to solve the robust model in (7), addressing the computational challenges posed by the decision-dependent nature of the uncertainties.
III-A Overview of Solution Methods
The presence of decision variables, which are solutions to the outer minimization problem, in the uncertainty set $\mathcal{D}_{2}$ poses significant complications to solving RO-DDU. Specifically, it changes the size and form of the set, making it difficult to reformulate robust problems with DDU sets in a tractable fashion. Indeed, the general RO-DDU problem is NP-complete, which can be proved based on the 3-Satisfiability problem (i.e., 3-SAT) [19]. Please refer to Appendix -A for the NP-Complete proof. Moreover, the uncertain parameter $d_{i,j}$ appears in both the objective function and the left-hand side of (7c), and the problem is a nonlinear program containing bilinear terms, such as $d_{i,j}x_{i,j}$, that cannot be solved directly.
To address these challenges, we propose the following two solution approaches. The high-level concepts of the two algorithms are illustrated in Fig. 2.
  • Robust Solution (RDDU): This approach combines linear programming (LP) duality, binary expansion, and linearization methods to reformulate the RO-DDU model into a large-scale MILP that can be solved by standard solvers.
  • Enhanced Robust Solution (e-RDDU): This approach introduces auxiliary variables and employs a clever transformation technique to reduce the problem size by eliminating redundant constraints during reformulation, thereby substantially improving computational efficiency.
Figure 2: Flowchart of solution approaches for the RO-DDU model. Two algorithms are shown: RDDU and e-RDDU.
III-B Robust Solution
The main challenge in solving the RO-DDU problem stems from the fact that the objective function (7a) and the robust constraints (7c) contain bilinear terms due to the inclusion of decision variables in the DDU set $\mathcal{D}_{2}$. When the uncertainty set is independent of the decision variables, standard techniques can be employed to convert the problem into an equivalent MILP robust counterpart using LP duality [15]. Unfortunately, $\mathcal{D}_{2}$ is decision-dependent, rendering our problem more complicated. To this end, we introduce a series of transformations to convert RO-DDU into an equivalent and solvable MILP form by transforming (7a) and (7c) into a set of linear equations.
@@ III-B Robust Solution
By using the preceding linearization steps, we obtain the following linear constraints that are equivalent to (11a)–(11c).
@@
The last term in (III-B) can be written explicitly as:
@@
Finally, the RO-DDU problem (7) can be reformulated as the following MILP, which can be solved by off-the-shelf solvers.
(RDDU) $\min_{{\bm{t}},{\bm{x}},{\bm{w}}}\quad\eta$ (21a)
@@
III-C Enhanced Robust Solution
Although the MILP reformulation (21) can solve the decision-dependent robust problem (RO-DDU) in (7), it suffers from weak relaxations due to the big-M terms in (12), leading to poor numerical performance. Furthermore, the size of the reformulation RDDU increases rapidly as the problem size, especially the upper bound of $x_{i,j}$, increases. To this end, we propose an alternative reformulation that can substantially improve the computational time. Specifically, we introduce new auxiliary variables and employ an alternative but equivalent form of $\mathcal{D}_{2}$ to reformulate the original RO-DDU problem. We first rewrite the constraints on $g_{i,j}$ in $\mathcal{D}_{2}$ as follows:
@@ III-C Enhanced Robust Solution
$\mathcal{D}_{2}(\hat{d},\bar{d},t,x,\Gamma_{2}):=\big{\{}d:~{}g\in\mathcal{G}^{\prime},$ (23)
$\!d_{i,j}\!=\!\bar{d}_{i,j}\!+\!\hat{d}_{i,j}\big{(}g_{i,j}+u_{i,j}x_{i,j}\big{)},\forall i,j\big{\}},$
@@
It can be observed that the key difference between the enhanced robust solution and the robust solution in Section III-B lies in the introduction of new binary variables ($v_{i,j}^{r}=1-t_{i,j}^{r}$) and replacing the upper bound of $g$ in (II-C) with (22). Now, the robust delay constraint (7c) can be expressed as:
@@
The last term in (III-C) can be given explicitly as:
@@
Overall, the RO-DDU problem in (7) can be reformulated as the following equivalent MILP:
(e-RDDU) $\min_{{\bm{t}},{\bm{x}},{\bm{w}}}\qquad\eta$ (31a)
@@
III-D Equivalence of RDDU and e-RDDU
@@
$\mathcal{U}({\bm{z}}):=\big{\{}\bm{\zeta}|\bm{A}\bm{\zeta}\leq{\bm{v}}+\bm{\psi}{\bm{z}}\big{\}},$ (32)
@@
It is easy to see that the RDDU reformulation relies on the result of Theorem III.1, while e-RDDU exploits Theorem III.2 to reduce the number of Big-M constraints. Therefore, we can infer that RDDU and e-RDDU are equivalent. Furthermore, the computational advantages of e-RDDU over RDDU stem from the fact that (35) has considerably fewer integer constraints compared to (34).
III-E Comparison Between RDDU and e-RDDU
@@
Remark III.3 (Computational Complexity).
The proposed RO-DDU model is a complex mixed-integer nonlinear program, proven to be NP-complete in Appendix -A, making it challenging to solve in its original form. The RDDU and e-RDDU reformulations transform the original problem into an MILP, solvable by off-the-shelf solvers. Consequently, the overall complexity of solving RO-DDU remains dependent on the solver’s choice and its computational capability. MILP itself is a challenging class of optimization problems, being NP-hard, and when treated as a decision problem, it is NP-complete [22]. Most MILP solvers incorporate heuristic algorithms during a pre-solve phase to effectively reduce the problem size before employing techniques such as branch-and-bound and branch-and-cut [22]. However, the computational time for these methods is primarily influenced by the problem size, particularly the number of integer variables, and the worst-case complexity remains exponential.
@@ IV Numerical Results
This section evaluates the performance of our proposed models through simulations, including comparisons with benchmarks and sensitivity analyses to examine the impact of key system parameters on the solution and system performance.
@@ IV-A Simulation Setting
We consider an EC system comprising $10$ areas and $10$ ENs (i.e., $I$ = $J$ = $10$), with larger networks explored in our sensitivity analyses. The edge network topology is based on the cities and locations of randomly selected Equinix edge data centers [^3] https://www.equinix.com/data-centers/americas-colocation, as similarly done in related work [23]. The network delay parameters, including the minimum delay $\bar{d}_{i,j}$ and maximum delay deviation $\hat{d}_{i,j}$, are generated using the global ping dataset444 https://wondernetwork.com/pings. Resource demands $\lambda_{i}$’s are generated using the Materna data trace [^5] http://gwa.ewi.tudelft.nl/datasets/gwa-t-13-materna, which includes performance metrics from distributed data centers, such as CPU cores, memory usage, and disk throughput over three months. This facilitates the simulation of realistic resource demands, with values ranging between $40$ and $60$ vCPUs, randomly generated based on a uniform distribution. The resource capacities of the ENs are randomly selected based on specifications of EC2 instances [^6] https://instances.vantage.sh.. Each EN consists of several EC2 instances. The unmet demand penalty parameters $s_{i}$ are generated using a uniform distribution over the interval $[40,50]$, which captures the variability in penalties that operators might incur depending on the criticality of service requirements and user expectations.
We consider three hardening levels for each link (i.e., $R=3$). The link hardening cost parameters (i.e., $h_{i,j}^{r},\forall i,j,r$) are generated as follows. Recall that $\Delta h$ is denoted by the incremental link hardening cost between two adjacent levels for each link, which is set to $0.2$ in our simulation. The first-level hardening costs $h_{i,j}^{\sf 1}$ are randomly generated from a uniform distribution within the interval $[1,1.05]$. The second and third-level link hardening costs are obtained by adding the incremental cost to the first-level hardening costs, i.e., $h_{i,j}^{\sf 2}=h_{i,j}^{\sf 1}+\Delta h$ and $h_{i,j}^{\sf 2}=h_{i,j}^{\sf 1}+2\Delta h,~{}\forall i,j$. These incremental costs are set to reflect realistic trade-offs between cost and performance improvement, consistent with established network hardening practices [12, 11]. The parameters $\gamma_{i,j}^{r}$, which represent the impact of the hardening decision on the link delay, are generated using a similar method. Specifically, we set $\gamma_{i,j}^{1},\forall i,j$ to $0.1$ and define $\Delta\gamma$ as an incremental impact factor (IF) between adjacent levels, which is set to $0.4$. Then, we have $\gamma^{2}=\gamma^{1}+\Delta\gamma$ and $\gamma^{3}=\gamma^{1}+2\Delta\gamma$. Thus, the IFs for all levels of hardening can be determined using only the IF for first-level hardening decisions ($\gamma_{i,j}^{r=1}$) and the incremental IF ($\Delta\gamma$), which simplifies conducting sensitivity analysis. It is important to note that the IF $\gamma_{i,j}^{r}$ falls within the range $[0,1]$, meaning the highest level, $\gamma_{i,j}^{r=3}$, must be less than $1$. To simplify, we assign distinct IFs within this range for the three hardening levels. The impact of the allocated workload is assumed to be uniform across all links, with the base case setting $u_{i,j}={\bm{u}}=0.1,\forall i,j$.
Parameter Value Parameter Value
$I,J$ $10$ $B$ 100
$R$ $3$ $\Gamma_{1}=\Gamma_{2}$ 15
$\gamma_{i,j}^{r=1}$ $0.1$ $h_{i,j}$ $U[1,1.05]$
$\Delta\gamma$ $0.4$ $\Delta h$ 0.2
$\rho$ $0.1$ $\Delta_{i}=\Delta$ 15
$\alpha$ $0.05$ $u_{i,j}$ $0.1$
TABLE III: System parameters
The system parameter values are detailed in Table III, which represents the default setting.During the sensitivity analysis, these key parameters will be varied to evaluate their impact on system performance. All the experiments are implemented in a Matlab environment using CVX [^7] https://cvxr.com/cvx/ and Gurobi 9.1.2 on a desktop with an Intel Core i7-11700KF and 32 GB of RAM. In CVX, we use the default precision setting, where the relative accuracy for the optimality conditions is approximately $10^{-6}$ to $10^{-8}$. For Gurobi solver, the default tolerance, i.e., MIPGap for MILP, is set to around $10^{-4}$ for mixed integer programs.
@@ IV-B1 Impacts of the link hardening costs and budget
As illustrated in Figs.3(a) and 3(b), increasing the hardening budget B leads to a decrease in the total cost since the platform can select more links for hardening, resulting in a reduction of the total cost. It is important to note that the delay constraints (1g) prioritize the closest ENs to meet the demand of each area, leading to only a subset of logical links being utilized. Clearly, the platform should only harden links with traffic traversing through them. Thus, as shown in Fig. 3(a), the cost becomes saturated after a certain budget value since the platform has already chosen the most critical links for hardening and has no incentive to harden more links even if the budget is redundant.
@@ IV-B2 Impacts of the uncertainty set
Figs. 4(a)-4(b) exhibit the effects of the uncertainty set on the optimal solution. It is evident from Fig. 4(a) that the total cost initially increases with the increment of $\Gamma$, resulting in a larger uncertainty set and a more robust solution. However, the total cost reaches saturation after a certain value of $\Gamma$, as the platform has already chosen an optimal set of links for hardening. Note that only a subset of ENs can serve demand from a specific area, causing certain logical links to remain unused for data traffic delivery. Therefore, the platform has no incentive to harden these links. Recall that $\bf{u}$ represents the influence of workload on link delay, as specified in $\mathcal{D}_{2}$. Fig. 4(b) suggests that a higher value of $\bf{u}$ corresponds to a greater link delay variation, leading to a higher total cost. These results are intuitive, as a larger uncertainty set yields a more robust solution, which, in turn, raises the cost.
@@
Figs. 5(a)-5(d) demonstrate the impacts of link hardening cost parameters (i.e., $\Psi$) and link hardening impact factors (i.e., $\gamma$) on the optimal objective and hardening cost (i.e., the total payment for link hardening). The incremental impact ($\Delta\gamma$) between two adjacent link hardening levels is also considered, where a higher value of $\Delta\gamma$ indicates a greater impact of level-2 and level-3 hardening, thereby offering more incentives for the platform to select links for level-2 and level-3 hardening. From $\mathcal{D}_{2}$, an increase in the link hardening impact results in lower link delays, leading to a reduction in overall delay cost. As $\Delta\gamma$ increases, so does its impact.
Fig. 5(a) shows that as $\Delta\gamma$ increases, the total cost decreases due to the significant reduction in link delay achieved through link hardening. This finding is further supported by Figs. 5(e)5(f). In addition, Fig. 5(b) indicates that the link hardening payment initially increases with the link hardening cost scaling factor $\Psi$ and then decreases after a certain value of $\Psi$, beyond which the hardening cost outweighs the benefit.
In Figs. 5(c)5(d), we set the incremental impact factor $\Delta\gamma$ to 0.3 and vary first-level impact factor ($\bm{\gamma}^{\sf 1}=\gamma_{i,j}^{\sf 1},\forall i,j$) from $0.1$ to $0.4$. A larger value of $\gamma^{\sf 1}$ leads to a lower link delay. Fig. 5(c) reveals that the total cost decreases as $\gamma^{\sf 1}$ increases, while Fig. 5(d) implies the platform invests more in link hardening as $\gamma^{\sf 1}$ increases. Finally, we vary $\Delta\gamma$ and the uncertainty budget $\Gamma$ to study their impact on the total cost and the link hardening payment in Figs. 5(e)5(f). It can be seen that an increase in $\Gamma$ leads to a higher cost due to the larger uncertainty set. Moreover, Fig. 5(f) suggests that the platform is willing to invest more money in link hardening as $\Gamma$ increases, indicating a higher level of uncertainty.